A 3.00 kg pumpkin hit the ground with 49.0 Joules of kinetic energy. How fast was the pumpkin falling? How far did the pumpkin fall?
1
Expert's answer
2020-11-23T10:28:57-0500
"K=0.5mv^2\\49=0.5(3)v^2\\v=5.72\frac{m}{s}"

"h=\frac{K}{mg}=\frac{49}{3\cdot9.8}=1.67\ m"